OpenStudy (anonymous):

Which expresses 6y : 9y in its simplest form? A.2 : 3 B.3 : 2 C.9 : 6 D.6 : 9

OpenStudy (anonymous):

i say that it is A

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

6/3=2 9/3=3 so its 2:3

OpenStudy (anonymous):

thank u love

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

anymore questions I can help with?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if not, can i get a medal?

OpenStudy (anonymous):

ok, go ahead and post it :D

OpenStudy (anonymous):

What is the ratio of x to y? 5x = 6y

OpenStudy (anonymous):

what are your answer choices? if you have none, thats fine i can still help :)

OpenStudy (anonymous):

The answer is 6/5

OpenStudy (anonymous):

A.6/5 B.5/6 C.1/5 D.1/6

OpenStudy (anonymous):

oh...I already posted the answer XD

OpenStudy (anonymous):

The answer is 6/5

OpenStudy (anonymous):

THANK YOOOOUUUUUU!!!!!!!!

OpenStudy (anonymous):

your welcome :)

OpenStudy (anonymous):

im here to help others :) thats what I do :D

OpenStudy (anonymous):

hehe ur cool

OpenStudy (anonymous):

On a soccer team, 11 out of 17 players surveyed say they had two or more siblings. The league has 850 players. Which is the best prediction of the number of players in the league that have two or more siblings? A.550 B.300 C.180 D.50

OpenStudy (anonymous):

thanks, not many people think im cool XD

OpenStudy (anonymous):

11/17 = 64.7% of kids have 2+ siblings of 850, 64.7% of kids or 549.95 (or 550) kids have 2+ siblings. so basically im saying the answer is 550 :)

OpenStudy (anonymous):

yayyy i got it right ok i got one more

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

nvm thats it love

OpenStudy (anonymous):

oh :) ok then, glad to help :D

OpenStudy (anonymous):

hehe ok see ya

OpenStudy (anonymous):

bye :D